OpenStudy (hellokitty17):

One solution of the equation 2x2 - x - 10 = 0 is five over two. What is the other solution? Numerical Answers Expected!

OpenStudy (welshfella):

factor the expression

OpenStudy (welshfella):

can you do that?

OpenStudy (welshfella):

2x2 - x - 10 = 2x^2 + 4x - 5x - 10

OpenStudy (welshfella):

factor the above by grouping

oregonduck (oregonduck):

If you start with 2x2 − x − 10 = 0 and divide everything by 2 to make the leading coefficient 1, you get x^2-(1/2)x-5 = 0 Now it turns out that for something like x^2+bx+c = 0, the two roots/solutions p and q multiply to the last coefficient c. In other words, pq = c Since one root is 5/2, we know that (5/2)q = -5 Now solve for q to get q = -10/5 = -2 So the other root is -2

OpenStudy (ali2x2):

lol

OpenStudy (welshfella):

someone nearly always gives the answer when we are trying to get the questioner involved so they can learn something...

oregonduck (oregonduck):

they do i explained it

OpenStudy (welshfella):

they would have learned a bit more if you hadn't interfered